OpenStudy (anonymous):

Please Help With This Problem! Kristi wants to invest $18,999 that she won in a raffle lottery. Reliable Bank is offering a 4-year CD with a 3.5% interest rate compounded annually. A- If Kristi wants to earn more than $2750 in total interest, can she earn this much with the CD from Reliable Bank? Explain. B- How much will Kristi have in her account after 6 years, if the bank offers a 6-year CD at the same rate? Explain.

OpenStudy (wolf1728):

Total = Principal*(1 + rate)^years Total = 18,999 * (1.035)^4 That equals 21801.7894888744 OR 21,801.79 rounded Interest = 21,801.79 -18,999 = 2,802.79 So, yes she can earn more than $2,700.00 in interest. Now, for part B

OpenStudy (wolf1728):

For 6 years: Total = Principal*(1 + rate)^years Total = 18,999 * (1.035)^6 Total = 18,999 * 1.2292553263 Total = 23354.6219452194 Equals 23,354.62 (rounded)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (wolf1728):

You are welcome :-)